MC VI: An electron in the Bohr model is in its first excited state (n of the electron 2), what is the linear momentum Q 11 0 12 1.0e-24 2.0e-24 J/s 3.0e-24 H/s 4.0e-24 5.0e-24 N/s rad/s kg m/s
Answer #1

p = me^2 / 2eo n h

where m = mass of the electron

e= charge of the electron

n = orbit , h = planck's constant

p = 9.11*10^-31*(1.6*10^-19)^2 / 2*8.85*10^-12*2*6.626*10^-34

= 1*10^-24 kg m/s

so 11)A 12) D is the correct answer